Exercise 18-09 Your answer is partially correct. Try again. Lendell Company has these comparative balance sheet data: Lendell Company Balance Sheets December 31 2020 2019 Cash $ 15,000 $30,000 Accounts receivable (net) 71,000 60,000 50,000 197,000 175,000 $344,000 $315,000 $50,000 $59,200 Mortgage payable (15%) 108,000 108,000 Common stock, $10 par 138,500 124,500 23,300 $344,000 $315,000 Inventory Plant assets (net) 61,000 Accounts payable Retained earnings 47,500

Additional information for 2020: 1. Net income was $25,700 2. Sales on account were $412,500. Sales returns and allowances amounted to $19,500. 3. Cost of goods sold was $199,800. 4. Net cash provided by operating activities was $45,600. 5. Capital expenditures were $23,800, and cash dividends were $10,100. Compute the following ratlos at December 31, 2020. (Round current ratio to 2 decimal places, e.g. 1.67. Round Accounts receivable turnover, Average colfethon p trloeu haventony turnover and Days in nventcry to raecinal Round Accounts receivable turnover, Average (a) Current ratio (b) Accounts receivable turnover (c) Average collection period (d) Inventory turnover (e) Days in inventory (f) Free cash flow 2.94 :1 6 times 122.0 days 3.6 times 101.3 days 8900

Help me solve for the red highlighted answers
Answer #1

1). Average Collection Period means how long time it takes to realise cash on credit sales.

Average collection period = Average accounts receivable balance for the year / Total net sales during the Year

Average.accounts.receivable.balance.for.the.year = \frac{71,000+60,000 }{2}

= $ 65,500

Total net sales during the Year = 412,500 - 19,500

= $ 393,000

Average collection period = (65,500/393,000)*365

= 60.83 days

2). Free cash flow

it is cash remain with the company after it pays its operational and capital expenditure

Free cash flow = operating cash flow - capital expenditures

= $ 45,600 - $ 23,800

= $ 21,800